1220) y=Ax+Dx^B is the particular solution of the first-order homogeneous DEQ: (x-y) = 2xy'. Determine A,B, & D given the boundary conditions: x=7 and y=5. Include a manual solution in your portfolio. ans:3
Answer #1

2ady = x-g o dut 2x ty + y = x dinding the whole equation by 24 Comparing it with differential equation of first order - InteDate Date II Page No. ухми 2- 36 2 = 4 см - се х22 23 с. 1 2 + c 2) U II, с = 1.css/ solution is yo x + 7.0.55 xth Comparing